OpenStudy (anonymous):

What is the area of the triangle? A. 42.14 cm² B. 35.28 cm² C. 21.07 cm² D. 17.64 cm²

OpenStudy (anonymous):

do you have a picture of it?

OpenStudy (anonymous):

okay give me a minute

OpenStudy (anonymous):

b

OpenStudy (anonymous):

17.64

OpenStudy (anonymous):

Its D

OpenStudy (anonymous):

ok can you answer more

OpenStudy (anonymous):

check the answer here

OpenStudy (anonymous):

Yeah ill check more

OpenStudy (anonymous):

The county animal park is shaped like a triangle. Two perpendicular sides of the park measure 2.3 miles and 3.4 miles. What is the area of the animal park? A. 3.91 mi² B. 5.7 mi² C. 7.82 mi² D. 11.4 mi²

OpenStudy (anonymous):

i would say D too.

OpenStudy (anonymous):

the formula for the area of a triangle is A= 1/2 bh so the area= 1/2 (base times height) A= 1/2 (9.8*3.6) A= 1/2 (35.28) A= 17.64

OpenStudy (anonymous):

A

OpenStudy (anonymous):

What is the area of a triangle with a base of 4.6 cm and a height of 7.8 cm? A. 12.4 cm² B. 17.94 cm² C. 24.8 cm² D. 35.88 cm²

OpenStudy (anonymous):

b

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Find the area of the triangle. A. 49 m² B. 245 m² C. 490 m² D. 980 m² http://static.k12.com/calms_media/media/1388500_1389000/1388522/1/13ad265b6f98390dcd08426fa462f29b4e086e2e/62361_FGA_U3_L6_V3_P.jpg

OpenStudy (anonymous):

b, 245